### v2

# Hierarchical and Modular Network on Non-prehensile Manipulation in General Environments

Abstract
For robots to operate in general environments like households, they must be able to perform non-prehensile manipulation actions such as toppling and rolling to manipulate ungraspable objects. However, prior works on non-prehensile manipulation cannot yet generalize across environments with diverse geometries. The main challenge lies in adapting to varying environmental constraints: within a cabinet, the robot must avoid walls and ceilings; to lift objects to the top of a step, the robot must account for the step’s pose and extent. While deep reinforcement learning (RL) has demonstrated impressive success in non-prehensile manipulation, accounting for such variability presents a challenge for the generalist policy, as it must learn diverse strategies for each new combination of constraints. To address this, we propose a modular architecture that uses different combinations of reusable modules based on task requirements. To capture the geometric variability in environments, we extend the contact-based object representation from CORN [17] to environment geometries, and propose a procedural algorithm for generating diverse environments to train our agent. Taken together, the resulting policy can zero-shot transfer to novel real-world environments despite training entirely within a simulator. We additionally release a simulation-based benchmark featuring nine digital twins of real-world scenes with 353 objects to facilitate non-prehensile manipulation research in realistic domains. Code, videos, and simulation benchmarks are available on the project website.
Figure 1: Illustration of our real-world domains. Our approach generalizes to diverse and unseen objects and environments. The object outlined in red represents the goal pose. Note how the environment geometry limits object and robot motions: in (a), the robot must reach and maneuver the object while avoiding the ceiling; in (b) and (c), the robot must circumvent walls to access the object; in (d), the robot must move the object over a barrier; in (e), the robot must first traverse the sink to contact under the crab, lift it across the sink wall, then reorient the crab to face forward; and in (i), the robot must consider its kinematics while manipulating the object on a high shelf. Each object and its pose are chosen so that it cannot be simply pick-and-placed.
11footnotetext: Equal Contribution.
I Introduction
Despite recent advances in robot manipulation, the practical deployment of robots in everyday environments like households remains challenging. One key reason is the robot’s inability to manipulate ungraspable objects. While much of prior work on manipulation centers around prehensile manipulation [57, 107, 27, 104], such approaches fall short in unstructured environments where objects are often ungraspable due to their geometry and the surrounding scene. To overcome this, robots must embrace non-prehensile manipulation, such as pushing, toppling, and rolling [44, 53, 17].
Recently, reinforcement learning (RL)-based approaches have achieved several successes in non-prehensile manipulation [45, 122, 17, 109]. However, these works have been limited to fixed objects in fixed scenes [45], general objects on flat tabletops [122, 17], or minor variations in objects and environments [109]. As such, no prior work addresses non-prehensile manipulation for novel objects and environments with arbitrary geometries as in Figure 1. Based on this observation, our objective is to extend RL to enable non-prehensile manipulation in such diverse setups.
The key challenge here lies in training a policy that can adapt to the constraints imposed by the given scene. For instance, consider the scenarios in Figure 1: for each domain, the robot is presented with a unique set of constraints, which may also evolve during an episode (Figure 1e). This requires the policy to not only model multiple distinct behaviors, but also rapidly switch between behaviors in response to minor changes in state. For example, in the scenario depicted in Figure 1e, as soon as the toy crab is positioned above the sink, the robot must quickly switch from a lifting skill to a translation skill. However, standard networks struggle to learn such high-frequency functions, a phenomenon known as spectral bias [99, 4, 88].
Figure 2: Illustration of computational structure for biological motor control (left) and our architecture (right). Each row compares analogous components in (1) acquiring sensory observations, (2) determining module activations, (3) modules representing groups of co-activated neurons, and (4) composing the modules to construct specific behaviors. The bar graphs denote the activation weight of each module.
Human brains, on the other hand, are extremely adaptive, and their computational structure for actions differs significantly from that of standard artificial neural networks. The computation of our brain is organized modularly [18], where a motor cortex orchestrates neural activities at the level of motor modules, a group of co-activating motor neurons [^1] Neurons connected to muscle fibers that trigger muscular contractions. [103, 72], to produce an action based on the current task [73, 101, 43]. By invoking different sets of modules in response to the current context, the motor system produces disjoint behaviors such as reaching or grasping without interference [25]. This is illustrated in Figure 2, left.
Inspired by this, we propose a modular and hierarchical policy architecture (Figure 2, right). In our architecture, the modulation network assumes the role of the motor cortex which determines the activations of modules, each representing a group of co-activated network parameters. The weighted combinations of these modules are then used to define the parameters of a base network, which has a fixed architecture but the parameters are determined online by the modulation network based on the current context, such as the environmental constraints. Unlike standard neural networks, this computational structure enables a single base network to model multiple functions, allowing the policy to produce distinct behaviors in response to even subtle changes in context.
Figure 3: Illustration of how distinct scenes map to distinct module activations that yield distinct behaviors, for a 5-layer network with 4 modules. The top-right colormap shows the activation of a particular module (column) for a particular layer (row): e.g., column 2, row 1 shows the activation of module 2 for layer 1. Opacity indicates the strength of module activation. The red object denotes the current object pose, and the blue object indicates the goal object pose. Ceil and NoCeil indicate the existence of a ceiling.
Figure 4: Success rates per architecture by parameter counts for a monolithic architecture (MLP and transformer) (x-axis, top) and number of modules for HAMnet (x-axis, bottom)
We refer to our modular policy architecture as Hierarchical and Modular Network (HAMnet). We discovered that, when we train the policy with HAMnet, it autonomously discovers modules and their activations that correspond to distinct manipulation strategies, such as reaching, lifting, or reorienting objects (Figure 3). Furthermore, as shown in Figure 4, we found that HAMnet scales much more gracefully with the number of parameters compared to a standard neural network in our simulated domains.
Another challenge in generalizing across diverse environments and objects is acquiring geometric representations from high-dimensional point cloud observations. One option is to use end-to-end training, yet jointly learning to encode point clouds incurs significant memory footprint, limiting its use with large-scale GPU-based simulators [58], which have become an essential tool for training RL policies for robotics. Pre-training representation models can mitigate this issue, but their effectiveness depends on the pretext task capturing features relevant to manipulation. Consequently, large off-the-shelf models are often unsuitable, not only from their high computational cost but also due to spurious geometric features that are irrelevant for contact-rich manipulation [119, 17].
Instead, we build on CORN [17], which learns object representations tailored for manipulation based on a pretext task that predicts the presence and location of contact between a gripper and object. The key insight is for contact-rich manipulation, it is important to capture what forces and torques can be applied to an object in the given state, which in turn depends on the presence and location of the contact between the object and the gripper. While CORN restricts itself to gripper-object contacts, we also need object-environment contacts in our problems as we generalize over environments. So, we introduce Universal CORN (UniCORN) that generalizes to contact affordances between two arbitrary geometries, based on a Siamese pre-training pipeline where a single encoder learns the representations for both an object and environment, and a decoder predicts the contacts between each environment point cloud patch and the object.
Our last contribution concerns domain design. To obtain a policy that generalizes to diverse environments, training environments must encompass a range of geometric constraints, while affording fast simulation for practical training. This is a non-trivial problem: unlike objects, environment assets for simulated RL training are not widely available, and manually designing them is costly. Therefore, we propose a procedural generation algorithm for constructing environments based on cuboid primitives. Online rearrangement of cuboids at different poses and dimensions yields wide coverage of geometric features that exist in the real world, such as walls, ceilings, slopes, and bumps. Further, the convex geometry of cuboids affords efficient dynamics simulation. Figure 3 shows example environments from our environment generation algorithm.
We show that by leveraging our framework, we can train a non-prehensile manipulation policy that can operate in diverse and novel environments and objects in a data- and time-efficient manner. We train our policy entirely in simulation and zero-shot transfer to unseen environments and objects in the real world. Furthermore, we provide a simulation-based benchmark comprising 9 digital twins of real-world environments and 353 objects to serve as a benchmark for non-prehensile manipulation for general environments and objects.
II Related Work
II-A Nonprehensile Manipulation
II-A1 Planning-based approaches
Prior works on planning-based non-prehensile manipulation use gradient-based optimization [67, 85, 68], graph-based search [56, 55, 66, 15, 49, 14], or a hybrid of both [10, 75]. To address the discontinuous dynamics arising from contact mode transitions, optimization-based works employ soft contact variables [67] or complementarity constraints [85, 68]. However, due to the imprecision of smoothed contact and the difficulty of precise constraint satisfaction, the resulting motions are difficult to realize in the real world.
On the other hand, graph-based methods [14, 40] can handle discrete dynamics transitions by representing the problem with a graph, where nodes represent robot states and contact modes, and edges encode the motion during the transition. This enables these methods to output more physically realistic motions for real-world deployment [15, 49]. However, to make the search tractable, these works restrict the diversity of motions, assuming quasi-static motions [14, 40] or predefined primitives [124, 49], limiting them to tasks with simple motions and sparse contact-mode transitions.
Other works combine optimization and sampling to accelerate planning [10, 75], yet remain too slow for online use due to the cost of searching large hybrid spaces with discontinuous dynamics. Further, most planning methods require knowledge of system parameters like mass and friction, which are difficult to estimate in real-world scenarios with varying objects and environments, harming practical real-world deployment.
II-A2 Learning-based approaches
Recent works leverage reinforcement learning (RL) to bypass the limitations of traditional planners by learning a policy that maps actions directly from sensory inputs [52, 117, 81, 118, 31, 121, 45, 122, 17, 109]. While this circumvents the computational cost of planning or the requirement of full system parameters, most of these works suffer from limited generalization across object geometries, since the policy is only trained on a single object [52, 117, 81, 118, 31, 121, 45]. Recent works incorporate point-cloud inputs [122, 17] or employ contact retargeting [109] to facilitate generalization across diverse object shapes, but none of these approaches adequately addresses the problem of generalizing across diverse environments using the full action space of the robot, as summarized in Table I.
TABLE I: Comparison of generalization capabilities for non-prehensile manipulation.
Object generalization General action space Environment generalization
HACMAN [122] O X X
CORN [17] O O X
Wu et al. [109] $\triangle$ $\triangle$ $\triangle$
Ours O O O
To generalize across diverse objects, HACMan [122] predicts an object-centric affordance map on its point cloud. While sample-efficient, this restricts robot motion to a hand-designed poking primitive, limiting their applicability in diverse environments. CORN [17] learns a policy over the full joint space of the robot, and generalizes over objects with a contact-based object representation tailored for manipulation. However, this work remains limited to a fixed tabletop due to the lack of environment representation. Wu et al. [109] retargets contacts from human demonstrations to determine robot actions for novel scenes, but their approach is limited to scenes similar to the original demonstration, as the actions are restricted to predefined skill sequences based on a primitive library. Like CORN, we train an RL policy over the full robot joint space to manipulate objects of general geometry. However, our approach generalizes across environments by leveraging our extended contact-based representation, UniCORN, and a modular network architecture, HAMnet.
II-B Multi-task Neural Architectures
In multi-task learning, a single model learns to do a family of related tasks by leveraging task synergies for improved performance and training efficiency [9]. As a single model must distinguish multiple tasks, it additionally takes context variables (e.g., task IDs) as conditioning inputs. The simplest approach for multi-task learning uses a monolithic architecture, which incorporates context inputs simply by concatenating them with network inputs. However, this design suffers from interference among tasks, as the neurons must handle multiple functions, leading to performance degradation [115, 42].
Recently, context-adaptive architectures have been proposed, where a separate network $g_{\phi}$ determines the parameters of base network $f_{\theta}$ from context inputs. See, for example, Figure 5 (a) and (c) for a comparison of monolithic and context-adaptive architectures. This separation allows the neural network to define a different function for each context, which mitigates interference [42]. Our architecture, HAMnet, also falls into this category.
Representative context-adaptive architectures include conditional normalization [82, 80, 65, 7], hypernetworks [37, 42], and modular architectures [34, 112, 97]. In conditional normalization [82] (Figure 5b), $g_{\phi}$ takes the context variable $z$ as input and applies feature-wise scale and bias $\{\gamma,\beta\}=g_{\phi}(z)$ to the intermediate features of the base network as $y=\gamma\odot f_{\theta}(x)+\beta$ where $\odot$ denotes element-wise multiplication. While effective, the expressivity of these architectures is limited to affine transforms, restricting its capacity [89]. In hypernetworks [37] (Figure 5c), $g_{\phi}$ affords broader expressivity, as it generates the entire set of base network parameters, $\theta=g_{\phi}(z)$, but suffer from poor training stability due to its large decision space over densely interacting parameters [71].
HAMnet is an instance of modular architectures (Figure 5d), where $g_{\phi}$ only predicts sparse activation weights $w$ of modules to determine $\theta$. Specifically, in modular architectures, $M$ denotes the number of modules, each of which is a network parameter, and the parameters of the base network are effectively formed as a weighted combination of these modules, such as $\theta=\sum_{i=1}^{M}w_{i}\theta_{i}$.
[…]
However, prior works on modular architectures for control assume that $z$ is given by the user and is discrete, such as pre-defined task IDs [84, 97, 112, 39]. In contrast, to generalize to novel, real-world environments in our problem, we need to infer $z$ from sensory observations such as environment and object point clouds. To overcome this, we design geometric encoders that map high-dimensional sensory observations to $z$.
Our architecture is most similar to Soft Modularization (SM) [112], with three key differences. Unlike SM, which predicts connectivities between all pairs of modules between neighboring layers, we directly predict module activations, which simplifies computation and reduces the output dimensions from $M^{2}$ connections to $M$ module activations. We additionally improve the computational efficiency by predicting all module activation weights in parallel, instead of predicting in series while conditioning on the preceding layer’s module activations. Lastly, we incorporate a gating mechanism [37] to enhance expressivity and boost policy performance.
II-C Representation learning on point clouds
To accelerate RL training with high-dimensional sensory inputs, prior works use representational pre-training [114, 29] to bootstrap RL agents. Different pretext tasks have been proposed for this purpose, such as point completion [105], orientation and category estimation [12, 41], or contrastive learning [110]. Inspired by the advances in natural language processing [22, 87] and image analysis [38], recent works adopt self-supervised learning (SSL) on patch-based transformers [116, 11, 76, 120, 1] for point cloud representation learning. These works reconstruct unseen geometric patches via either autoregressive prediction [11] or masking [116, 76, 120, 1] to learn rich geometric representations, achieving state-of-the-art results in shape classification and segmentation [11].
Despite their success in general-purpose vision tasks, these representations are unsuitable for robot manipulation for two reasons. First, these models attempt to predict the missing patches in a point cloud, which forces the encoder to focus on encoding information about the object’s shape. However, knowing the exact shape of the object is often sufficient but unnecessary for manipulation. For example, manipulating a toy crab in Figure 1e does not require knowledge of the exact shape between its legs, as that area is tightly confined and cannot be contacted by the robot or the environment. Second, capturing such spurious details requires a large model, which degrades training efficiency [28, 94] and policy performance [119]. In contrast, we extend CORN [17] to pre-train a representation to encode contact affordances among arbitrary geometry pairs, shown to be effective for robot manipulation.
II-D Modularity in biological networks
Modularity in biological neural networks is a key principle underlying adaptation and learning [96]. In vertebrate motor systems, muscle synergies [103, 20, 72, 61, 23] serve as modules of movement that abstract muscle control, representing a coordinated contraction of a set of muscles to produce a desired behavior, such as the synchronized activation of the quadriceps and hamstrings for walking [23].
Modularizing motor control in this way provides several benefits. When adapting to a particular context, the central nervous system (CNS) can dictate behaviors using sparse, low-dimensional signals that activate specific muscle groups [5]. Compared to controlling individual motor neurons, this affords rapid switching between distinct motor skills like reaching and grasping depending on the context [73, 101]. Further, synergies can be reused across behaviors, producing diverse movements such as pinch- or power-grasps from a limited set of modules [86], which facilitates learning by recombining and adapting existing modules [18, 25]. In our work, we incorporate these principles to design our architecture.
II-E Skill discovery in RL
In skill discovery, Unsupervised RL (URL) aims to find reusable skills using task-agnostic objectives like state coverage [79, 51, 64, 83, 113, 8] or skill diversity [35, 26, 93, 47, 77]. However, without task-specific priors, such intrinsically motivated methods often fail to make meaningful interactions in high-dimensional domains without engineered bias [3, 102]. While METRA [78] was shown to scale to pixel-based tasks, it remains prone to degenerate behaviors, such as lying still in varied poses in humanoid control domains [102]. On the other hand, our framework learns task-specific skills grounded in the training domain, yielding interpretable skills that transfer to real-world robot manipulation.
II-F Mitigating Spectral Bias in RL
A central challenge in multi-task learning is enabling networks to adaptively switch functions under subtle shifts in its input. While our work focuses on the design of network architecture, one complementary line of work investigates improving input representations with high-frequency encodings, such as Learned Fourier Features (LFFs)[99], to enhance the network’s sensitivity to sharp transitions. While promising, the effectiveness of LFFs remains inconclusive in RL, with conflicting claims on the role of different frequency components [48, 111, 60]. Moreover, LFFs lead to an exponential increase in input dimensionality [6] and may degrade training stability from the increased variance, impairing generalization and robustness to noise [60]. These limitations suggest the need for an architecture-level solution, and not just the input representation. Still, one potentially promising future direction is to see whether high-frequency embeddings can complement the modular architecture by increasing network sensitivity to distinct contexts when determining module activations.
Figure 6: Overall method overview. Our framework consists of four main components: a modular policy trained with RL, contact-based representation pre-training, a procedural domain generation scheme for environment geometries, and a simulation-to-real transfer method for real-world deployment. Dashed blocks indicate external inputs.
[…]
III Methodology
We consider the non-prehensile manipulation problem, where a robot arm with a fixed base moves an object to a target pose in environments of general geometry, e.g., kitchen sinks, cabinets, and drawers (Figure 1). We model this problem as a Markov Decision Process (MDP), represented as a tuple $(S,A,P,r,\gamma)$ denoting state space $S$, action space $A$, state transition model $P(s_{t+1}|a_{t},s_{t})$, reward model $r(s_{t},a_{t},s_{t+1})$, and discount factor $\gamma$. Our objective is to obtain policy $\pi$ that maximizes the return $R_{t}=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}[\sum\gamma^{t}r(s_{t},a_{t},s_{t+1})]$ via a sequence of non-prehensile actions.
Figure 6 presents an overview of our framework. We leverage deep RL in a parallel GPU-based simulation [58] to train a modular policy (Section III-B) using the pre-trained point cloud representation (Section III-C) on procedurally generated domains (Section III-D). We distill the resulting policy for real-world deployment via teacher-student distillation (Appendix -D). All pre-training, policy training, and distillation stages happen entirely in a simulation.
III-A MDP Design
Our state space $S$ consists of robot joint state $x^{q}_{t}$, end-effector pose $x_{t}^{EE}$, object geometry $G_{o}$, environment geometry $G_{e}$, and goal pose $T_{g}$. In the simulation, the agent additionally receives physics parameters $\nu$ and object state $x^{o}_{t}$. We represent all poses as 3D translation and 6D orientation to facilitate learning [123], and the goal is given as a relative pose from the current object pose. Object and scene geometries are given as surface-sampled point clouds.
Our action space $A$ consists of joint residuals $\Delta q\in\mathbb{R}^{7}$ and controller gains, parameterized by proportional gains $k_{p}\in\mathbb{R}^{7}$ and the damping ratio $\rho\in\mathbb{R}^{7}$ that maps to the damping gain $k_{d}$ as $\rho\sqrt{k_{p}}$, following [59, 45]. The resulting torque for each joint is computed as $\tau=k_{p}\Delta q-k_{d}\dot{q}$. While prior works adopt Cartesian-space actions [122, 17], we adopt joint-space actions, which enables direct control of individual robot links to avoid collisions against the environment during manipulation.
The reward $r(s_{t},a_{t},s_{t+1})$ in our domain is defined as a sum of the task success reward $r_{s}$, goal-reaching reward $r_{r}$ and the contact-inducing reward $r_{c}$: $r=r_{s}+\lambda_{r}r_{r}+\lambda_{c}r_{c}$, where $\lambda_{r}$ and $\lambda_{c}$ are scaling coefficients for the respective rewards. Since $r_{s}=\mathds{1}_{suc}$ is sparsely given, we incorporate shaping rewards $r_{r}$ and $r_{c}$ as potential functions of the form $\gamma\phi(s^{\prime})-\phi(s)$ with the discount factor $\gamma\in[0,1)$, which preserves policy optimality [69]. Specifically, we have $\phi_{r}(s)=-\log{(c_{g}\cdot d_{o,g}(s)+1)}$ for $r_{r}$, and $\phi_{c}(s)=-\log{(c_{r}\cdot d_{h,o}(s)+1)}$ for $r_{c}$, where $c_{g},c_{r}\in\mathbb{R}$ are scaling coefficients for the distance-based potential functions; $d_{o,g}(s)$ is the relative distance between the current object and the goal pose, based on the bounding-box distance [2]; $d_{h,o}(s)$ is the hand-object distance between the object and the tip of the end-effector. Task success is achieved when the object’s pose is within 0.1m and 0.1 radians of the target pose. The episode terminates if (1) the object reaches the goal, (2) the object is dropped from the workspace, or (3) the episode reaches the timeout of 300 simulation steps. Table VII summarizes our MDP design, and details on reward coefficients are in Table VIII.
III-B HAMnet-based architecture
Our architecture, shown in Figure 7, consists of three main components: the geometry encoder (red), modulation network (green), and base network (blue). Our proposed modular architecture, HAMnet, consists of the modulation and base networks. Since we use PPO, an actor-critic algorithm, our base network outputs both value and action.
The geometry encoder processes three types of point cloud inputs: the global scene cloud, capturing the overall geometry of the scene; the local scene cloud, detailing the nearby scene that surrounds the object; and the object point cloud, representing its surface geometry. Each cloud is patchfied, tokenized, and embedded by the pre-trained geometry encoder (Section III-C), yielding latent geometric embeddings $z^{(G)}_{\text{1:N}}$, $z^{(L)}_{\text{1:N}}$, $z^{(O)}_{\text{1:N}}$, for global, local, and object embeddings, respectively. Details on point cloud acquisition are in Appendix -F.
[…]
We then use $w$ and $g$ to build the base network parameters $\theta$. For each layer $j\in[1\dots L]$, $w=\{w_{i,j}\}_{j=1}^{L}\in\mathbb{R}^{L\times M}$ act as M module-wise weighting coefficients, passed through softmax to ensure $\sum_{i=1}^{M}w_{i,j}=1$. The gating factor $g=\{{g_{j}}\}_{j=1}^{L}\in\mathbb{R}^{L\times D_{j}}$ is a feature-wise multiplier for each layer, with $D_{j}$ denoting the number of output dimensions of layer $j$. Together, $\theta$ is constructed as a weighted composition of modules followed by gating, such that $\theta=\{(\sum_{i=1}^{M}w_{i,j}\theta_{i,j})\odot g_{j}\}_{j=1}^{L}$.
[…]
Note that since the base network has both the actor and critic, we keep separate sets of modules for each of them, denoted $\{\theta_{i,j}^{(a)}\}_{i=1}^{M}$, and $\{\theta_{i,j}^{(v)}\}_{i=1}^{M}$ for layer $j$. Our module activation weights and gating factor also consist of weights for a value and action, $w=\{w^{(v)},w^{(a)}\}$ and $g=\{g^{(v)},g^{(a)}\}$. To make a prediction, the base network gets instantiated twice for actor and critic; in the former case, the network uses the weight $\theta^{(a)}_{j}=\sum_{i=1}^{M}w_{i,j}^{(a)}\theta_{i,j}\odot g^{(a)}_{j}\}_{j=1}^{L}$, and in the latter, the network uses $\theta^{(v)}_{j}=\sum_{i=1}^{M}w_{i,j}^{(v)}\theta_{i,j}\odot g^{(v)}_{j}\}_{j=1}^{L}$. These details are omitted in the figure for brevity.
Figure 8: Our pre-training architecture consists of a geometry encoder (red) and a contact decoder (green). The same geometry encoder operates on each point cloud A and B in a Siamese fashion to produce local patch embeddings $z^{A}_{1:N-1}$, $z^{B}_{1:N-1}$ and global embeddings $z^{A}_{N}$, $z^{B}_{N}$. The contact decoder (green) predicts contact between each patch $z^{A}_{i}\in z^{A}_{1:N-1}$ and $z^{B}_{N}$. The bottom block details the procedure to patchify and tokenize point clouds.
III-C Training UniCORN
We design our representation pre-training task on estimating the presence and location of contact between two point clouds, A and B.
III-C1 Pre-training data generation
To acquire data for pre-training, we generate a dataset containing pairs of objects represented as point clouds, and contact labels indicating the presence and location of contact. Using the objects from DexGraspNet dataset [106], we generate the data by (1) sampling near-contact object configurations, (2) creating the point clouds by sampling points from the surface of each object, and (3) labeling contact points based on whether they fall within the other object. To account for possible scale variations between geometries, we sample the point clouds at varying densities and scales during this process. For details on the data generation pipeline, see Appendix -B.
III-C2 Network Architecture
Figure 8 shows our pretraining network architecture, comprising the geometry encoder and the contact decoder. The encoder takes the point clouds of objects A and B, denoted $x_{A},x_{B}$ as inputs, mapping the patch-wise tokens from $x_{A}$ and $x_{B}$ and a learnable [EMB] token to local patch embeddings $z_{1:N-1}^{A}$, $z_{1:N-1}^{B}$ and global embeddings $z_{N}^{A}$,$z_{N}^{B}$. Afterward, the decoder takes $(z_{i}^{A},z_{N}^{B})_{i=1}^{N}$ and predicts the presence of contact at each of $i$-th local patch of object A with object B. The overall network is trained via binary cross-entropy against the patch-wise contact labels. During training, we alternate the roles of A and B (i.e., A-B and B-A) to ensure that we also use the global embedding of A and predict the contact at a patch of B.
Figure 8 (bottom) shows the procedure to tokenize the point clouds. In line with previous patch-based transformer architectures for point clouds [76, 11, 17], we first patchify the point cloud by gathering neighboring points from representative center points. These center points are selected via farthest-point sampling (FPS), and the points comprising the patches are determined as the k-nearest neighbors (kNN) of the patch center. These patches are normalized by subtracting their center coordinates, and a small MLP-based tokenizer [17] embeds the shape of each patch. Afterward, we add sinusoidal positional embeddings of the patch centers to the patch tokens to restore the global position information of each patch.
III-D Procedural domain and curriculum generation
Figure 9: Our pipeline for environment generation composes different environmental factors, such as walls, ceilings, and plates at different elevations for each axis, to construct geometrically diverse environments.
To create diverse environments and support curriculum learning for training our policy, we develop a procedural generation scheme for constructing environments as a composition of cuboidal primitives. Since we construct environments by dynamically rearranging existing geometric entities, it integrates well with most GPU-based simulators [58, 33] that prohibit spawning new assets after initialization.
Our procedural pipeline, shown in Figure 9, comprises two main components: domain interior and exterior generation. The interior includes planar and interim plates arranged laterally and longitudinally, where planar plates form elevated surfaces and interim plates form sloped ramps. Their dimensions, elevations, and angles are randomly sampled to produce diverse topographies, yielding features like bumps, valleys, and steps (Figure 3). The exterior consists of walls and ceilings that impose accessibility constraints, where their presence, height, and ceiling type (nominal or tight) are randomly sampled. The proportion of ceiling types controls the difficulty of workspace accessibility, since the nominal ceiling is generated with sufficient clearance, whereas tight ceilings leave a narrow margin relative to the object’s height. Details on the procedural generation pipeline are in Appendix -C.
As our procedural generation pipeline is fully parameterized, the sampling distributions of environmental parameters can be dynamically adjusted during training. This enables curriculum learning, where task complexity is incrementally increased throughout training. Specifically, we employ a curriculum for robot initialization and ceiling types. To facilitate this, we additionally introduce two types of robot initializations: near and random. In the near configuration, the end-effector begins within a 0.1m radius of the object based on collision-free inverse kinematics solutions from CuRobo [98]; in the random configuration, a collision-free joint configuration is uniformly sampled within the robot’s joint limits. Early in training, we preferentially sample near initializations and nominal ceilings to encourage interaction with the object. As training progresses, we linearly increase the proportion of random initializations and tight ceilings, encouraging the policy to develop obstacle-aware maneuvers for approaching objects from arbitrary configurations.
IV Experimental Results
IV-A Overview
Our goal is to evaluate the following claims: (1) our modular architecture, HAMnet, affords data-efficient training for a policy that generalizes over large domain diversity, compared to monolithic or hypernetwork architectures; (2) our contact-based representation, UniCORN, affords data-efficient training for a robot manipulation policy in geometrically rich domains compared to an off-the-shelf self-supervised representation; (3) our framework affords real-world transfer and generalization to novel environment geometries despite only training in a simulator with synthetic environments.
TABLE II: Comparison between baselines regarding architecture and representation.
Model Name Model Architecture Representation
UniCORN-HAMnet (ours) HAMnet UniCORN
UniCORN-Hyper Hypernetwork UniCORN
UniCORN-SM Soft-Modularization [112] UniCORN
UniCORN-Transformer Transformer UniCORN
UniCORN-Mono MLP UniCORN
PointGPT-HAMnet HAMnet PointGPT [11]
E2E-HAMnet HAMnet End-to-end
To evaluate our claims, we compare the performance of our proposed model (UniCORN-HAMnet) with the baselines summarized in Table II. These baselines explore alternative choices in network architecture or representation. UniCORN-Hyper uses a hypernetwork [50, 91] to predict base network parameters. UniCORN-SM is a variant of a modular architecture using Soft Modularization [112]. We include two variants of standard monolithic architectures, UniCORN-Mono and UniCORN-Transformer, respectively using an MLP and a transformer. PointGPT-HAMnet and E2E-HAMnet considers alternative choices in the representation, where the former replaces the pretrained UniCORN with a PointGPT encoder [11], while the latter jointly trains the representation model end-to-end. All architectures are configured to have a similar number of trainable parameters up to the architectural constraints. Additional details on the baselines are in Appendix -G.
IV-B Simulation experiment
To train our policy, we use a Franka Research 3 (FR3) arm manipulating a subset of 323 objects from DexGraspNet dataset [106] on the procedurally generated environments as in Section III-D. We train each baseline using PPO [92] with identical hyper-parameters, spanning 2 billion environment interactions across 1024 parallel environments in Isaac Gym [58]. Detailed hyperparameters for policy training are described in Appendix -G. We consider two metrics: data efficiency and time efficiency.
Figure 10: Training progression. For each baseline, we show the mean (solid) and standard deviation (transparent) of success rates across three seeds. The interaction steps are aggregated across 1024 parallel environments.
To support our claim on training efficiency, we consider the training progression plot in Figure 10. Overall, modular architectures (UniCORN-HAMNet and UniCORN-SM) achieve the best data efficiency, with the mean success rates of 75.6% and 70.9% after training. In contrast, monolithic architectures show lower performance, regardless of whether conditioning is given by concatenation (UniCORN-Mono, 49.8%) or self-attention (UniCORN-Transformer, 43.9%). UniCORN-Hyper performs best among non-modular architectures at 56.5%, indicating the adaptivity of the network expedites policy training. UniCORN-SM (70.9%) further improves over hypernetworks, as its modularity affords reuse of network modules and reduces the learning complexity by predicting sparse module activations rather than parameters of individual neurons. Lastly, UniCORN-HAMNet (75.6%) outperforms UniCORN-SM from the increased expressivity of the gating mechanism. To evaluate the representational efficacy of UniCORN, we also compare UniCORN-HAMNet to E2E-HAMNet and PointGPT-HAMNet. End-to-end training (E2E-HAMNet) degrades performance (47.9%), emphasizing the utility of pre-training; while PointGPT-HAMNet performs better (66.0%), it still underperforms UniCORN-HAMNet due to the overhead from spurious geometric details and increased embedding dimensions.
Figure 11: Per-epoch training time comparison across all baselines, measured on identical hardware (NVIDIA A6000). Error bars represent two standard deviations.
Since all our training happens in simulation, training time is also an important factor. To evaluate the time-efficiency of each baseline, we measure the per-epoch training time in Figure 11. While the monolithic MLP architecture (UniCORN-Mono) is the fastest (1.12s) due to its simplicity, modular architectures (UniCORN-HAMNet and UniCORN-SM) follow closely at just 1.33s and 1.84s, which shows that determining module-level activation adds minimal overhead; between the two, UniCORN-HAMNet achieves faster training than UniCORN-SM from the streamlined prediction of module activations. In contrast, UniCORN-Hyper takes the longest (5.31s) due to the cost of predicting the full set of base network parameters. To contextualize the overhead, a standard transformer (UniCORN-Transformer) takes around 3.30s. We also compare with representational baselines: PointGPT-HAMNet and E2E-HAMNet. The large encoder of PointGPT-HAMNet significantly lags training, averaging 4.40s per epoch, about 3.3 times slower than UniCORN-HAMNet. While E2E-HAMNet uses the same encoder architecture as UniCORN-HAMNet, it suffers from 2.9$\times$ slower training due to the overhead of co-training the representation model.
IV-C Real world experiment
TABLE III: Results on 9 unseen real-world domains.
Domain Object Success rate Domain Object Success rate
Cabinet Bulldozer 4/5 Top of cabinet Bulldozer 3/5
Heart-Box 3/5 Crab 4/5
Sink Bulldozer 5/5 Basket Bulldozer 3/5
Angled Cup 4/5 Heart-Box 5/5
Drawer Bulldozer 4/5 Grill Bulldozer 5/5
Pencil case 3/5 Dino 4/5
Circular bin Bulldozer 4/5 Flat Bulldozer 5/5
Pineapple 3/5 Nutella 3/5
Suitcase Bulldozer 4/5 Total 78.9%
Candy Jar 5/5
To validate the real-world applicability and generalizability of our framework, we evaluate our policy in 9 real-world domains with novel everyday scenes and objects (Figure 1). We test two objects in each domain: one object (a toy bulldozer), shared across all domains, and one random object (Figure 12(b)), each with five trials at different initial and goal poses.
For each scene, we mount four RealSense D435 cameras to observe the point clouds from multiple viewpoints, ensuring sufficient visibility of the object during execution (Figure 12(a)). To distinguish the object cloud from the environment cloud, we use SAM [46] to designate the initial object segmentation mask and utilize Cutie [13] to track the object during manipulation. We use FoundationPose [108] to estimate the object’s relative pose from the goal pose, using the view with the best visibility of the object (largest object segmentation mask) among the four cameras. We generate the environment point cloud by combining and filtering the point clouds from the depth cameras. We replaced the robot’s gripper to accommodate narrow environments, wrapped with a high-friction glove to reduce slipping. Further details on the real-world setup are in Appendix -I.
Table III shows the results of our policy across 9 real-world domains. Overall, our policy demonstrates 78.9% success rate, indicating that our framework facilitates the policy to transfer to diverse, unseen real-world environments, despite only training in a simulation. The main failure modes of our policy are, in decreasing order of frequency: torque limit violation (5.56%); policy deadlock (4.44%); dropping objects (4.44%); getting blocked by the environment (3.33%); and perception error (3.33%). Detailed descriptions of these failure modes are in Appendix -I2.
(a) Example of our scene layout in the cabinet domain.
(b) Our real-world objects, all unseen during policy training.
Figure 12: Our real-world experimental setup.
IV-D Emergence of skills in HAMnet
Figure 13: (inner block) UMAP projection of the modulation embedding $z_{\tau}$, colored by clusters from HDBScan. Unclustered points are in black. (outer block) Isolated view of each cluster, colored by a representative domain parameter, such as ceiling height or goal direction (to the left or right). The camera icon per each box denotes the viewpoint. The rendered scene shows a domain and state that generated an embedding in a cluster, with the red and blue objects indicating the current object pose and goal pose respectively.
We show that HAMnet automatically discovers different manipulation skills and learns to sequence them. To do this, we inspect the modulation embedding $z_{\tau}$ (see Figure 7), which decides the activation weight of each module. We collect a dataset of $z_{\tau}$ by running a trained policy in 25,000 randomly sampled episodes in simulation. Since the high-dimensional $z_{\tau}$ is hard to interpret, we project $z_{\tau}$ into a three-dimensional manifold using UMAP [63] to visualize its structure.
To show that HAMnet discovers different skills, we apply HDBSCAN [62] to these projections of $z_{\tau}$. Figure 13 shows the result. The inner block of Figure 13 shows the extracted clusters with different colors. We find that, without any manually designed bias or knowledge, these clusters naturally emerge and have semantically interpretable behaviors, such as lifting (red), reaching (blue), reorienting with (yellow) and without a ceiling (bright yellow), translation with (purple) and without a ceiling (bright purple), and dropping objects (green).
[…]
(a) An episode where the agent lifts an object across a platform.
(b) An episode where the agent drops an object to a lower elevation.
Figure 14: Illustration of how our architecture learns to use different skills. Color bar in each subfigure shows the cluster labels of $z_{\tau}$ at each step, and the bottom shows the domain rendering of representative keyframes. The red object is the current object pose, and the blue object is the goal pose.
To check if HAMnet can use these skills in sequence, we analyze how $z_{\tau}$ changes throughout a task. To highlight the transitions, we label $z_{\tau}$ at each step of an episode based on the precomputed HDBScan clustering shown in Figure 13. Figure 14 shows that the agent switches between behavioral clusters based on its internal subgoal: in Figure 14(a), the robot initiates with a reaching skill to approach the object while avoiding obstacles. Afterward, the robot transitions to lift the object to the top platform. After a successful lift, the robot reorients the object to match the target orientation. Lastly, the robot translates the object to its target pose. The sequence of $z_{\tau}$ changes when the problem changes: when the robot has to drop an object to a lower platform instead (Figure 14(b)), the robot follows a different sequence (reach-drop-reorient-translate). This demonstrates that our architecture can (1) discover its own subgoals and (2) activate different modules to achieve different subgoals.
IV-E Simulated Benchmark in Realistic Domains
We release a simulated digital twin of our nine real-world setups as a benchmark for non-prehensile manipulation (Figure 15). The environment mesh is built using CAD, Nerfstudio [100], and Polycam. Our benchmark comprises 353 objects: 9 custom scans from the real world, 21 from GSO [24], and 323 from DGN [106]. For each domain and object pair, we sample 5 stable initial- and goal-poses and 128 random collision-free robot initializations to evaluate generality. Appendix -H4 details domain configurations and provides baseline results.
Figure 15: Sample environments in our simulated benchmark. From the top left: sink, circular bin, suitcase, basket, grill, and cabinet. The axis on each object indicates its current pose, while the other axis represents the target pose.
V Conclusion
In this work, we propose a novel framework for non-prehensile manipulation in general environments via deep reinforcement learning in a simulation. Our framework comprises a modular architecture (HAMnet), a contact-based object and environment representation (UniCORN), and a procedural domain generation algorithm for diverse environment geometries. Compared to conventional architectures and standard representations, our framework facilitates data- and time-efficient training of a policy that generalizes to diverse and unseen scenes. Despite solely training in synthetic environments in a simulation, our policy zero-shot generalizes to unseen real-world environments and objects. Overall, our combined framework achieves state-of-the-art performance in non-prehensile manipulation of general objects in general environments.
V-A Limitations
Despite promising results, our approach has several limitations that can be addressed in future work:
Improved efficiency of HAMnet. In HAMnet, the parameters of the base network are updated at every step. However, our qualitative analysis (Section IV-D) shows that $z_{\tau}$, and the resulting predictions for the module activation weights, remain stable for prolonged periods until a transition triggers a change, such as a successful lift. Thus, reusing the predicted network parameters over multiple steps can potentially reduce the computational overhead.
Dynamics-aware object representation. While UniCORN effectively represents object and environment geometries, it neglects dynamic properties like mass and inertia, which are critical for maneuvering objects with unusual dynamics, like roly-poly toys with non-uniform mass distributions. As such, one intriguing future research direction is to extend the representation to include dynamics information, potentially by incorporating memory [16, 36].
Generating fine-grained environment features. Our procedural generation pipeline relies on cuboidal primitives, limiting the diversity of fine-grained geometric features (e.g., textures, curvatures, small overhangs). While our experiments in the Grill, Drawer, and Circular Bin environments show that the policy can still adapt to uneven and curved surfaces, diversifying procedural generation through approaches like geometric generative models [95], may enhance the policy’s environmental generalization capability.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant and National Research Foundation of Korea (NRF) funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST)), (No.2022-0-00311, Development of Goal-Oriented Reinforcement Learning Techniques for Contact-Rich Robotic Manipulation of Everyday Objects), (No. 2022-0-00612, Geometric and Physical Commonsense Reasoning based Behavior Intelligence for Embodied AI), (No. RS-2024-00359085, Foundation model for learning-based humanoid robot that can understand and achieve language commands in unstructured human environments), (No. RS-2024-00509279, Global AI Frontier Lab).